0 0
Add a comment Improve this question Transcribed image text
Answer #1

apply power = eenrgy/time

time = 262500/(75

time = 3500 secs

t = 3500/60 = 58.33 mins

t = 0.97 Hrs
--------------------------------------------

1Psi =2.0364 in of Hg

so 140 psi = 140 *2.306

= 286 in of Hg